Simplify 3^1/2 times 3^1/2
Found 2 solutions by Boreal, stanbon:
Answer by Boreal(15235)

Multiply and add exponents
3^(1/2) * 3(1/2)=3^1=3

Answer by stanbon(75887)

Simplify 3^1/2 times 3^1/2
----
Ans: (3*2)^(1/2) = 6^1/2
-------------------
Cheers,
Stan H.
----------